OpenStudy (anonymous):

After 6 years, an investment is worth $8479. If the money was invested at 9% interest compounded semiannually, what was the initial investment?

OpenStudy (anonymous):

A=P(1+(r/n)^nt

OpenStudy (anonymous):

solve \[8479=P(1+\frac{.09}{2})^{12}\]for P

OpenStudy (anonymous):

\[\frac{8479}{(1.045)^{12}}=P\] now a calculator